OpenStudy (anonymous):

Need help with a few precalculus questions. Simplify the expression. quantity one minus cotangent of x divided by quantity tangent of x minus one

OpenStudy (anonymous):

cotx

OpenStudy (anonymous):

thank you. how about this one

OpenStudy (anonymous):

1-cotx/tanx-1 =( 1-1/tanx)/tanx-1 =( tanx-1)/(tanx*tanx-1) =1/tanx =cotx

OpenStudy (anonymous):

Verify the identity. quantity one minus sine of x divided by cosine of x equals cosine of x divided by quantity one plus sine of x

OpenStudy (anonymous):